# Fast constrained sampling in pre-trained diffusion models

### v2

@@ Abstract
Large denoising diffusion models, such as Stable Diffusion, have been trained on billions of image-caption pairs to perform text-conditioned image generation. As a byproduct of this training, these models have acquired general knowledge about image statistics, which can be useful for other inference tasks. However, when confronted with sampling an image under new constraints, e.g. generating the missing parts of an image, using large pre-trained text-to-image diffusion models is inefficient and often unreliable. Previous approaches either utilize backpropagation, making them significantly slower and more memory-demanding than text-to-image inference, or only enforce the constraint locally, failing to capture critical long-range correlations. In this work, we propose an algorithm that enables fast and high-quality generation under arbitrary constraints. We observe that, during inference, we can interchange between gradient updates computed on the noisy image and updates computed on the final, clean image. This allows us to employ a numerical approximation to expensive gradient computations, incurring significant speed-ups in inference. Our approach produces results that rival or surpass the state-of-the-art training-free inference approaches while requiring a fraction of the time. We demonstrate the effectiveness of our algorithm under both linear and non-linear constraints. An implementation is provided at this link.
Diffusion models, conditional sampling
@@ 1 Introduction
(Rout et al., 2023) (Yu et al., 2023) (He et al., 2024)
Figure 1: When tasked with completing the missing half of an image, previous methods are slow and fail to capture the important long-range dependencies between image pixels. The proposed algorithm generates a plausible image at a fraction of the time.
The development of large text-to-image models (Nichol et al., 2022; Rombach et al., 2022; Ramesh et al., 2022; Saharia et al., 2022b) has made denoising diffusion (Sohl-Dickstein et al., 2015; Ho et al., 2020) the go-to approach for capturing complex data distributions in high-dimensional spaces, such as images. By training on billions of text-image pairs, these models have acquired general knowledge about the image space, beyond text-to-image generation. This knowledge has been useful in quickly adapting to new conditions (Zhang et al., 2023) as well as utilizing model features to solve downstream image tasks (Tang et al., 2023).
The simplest way to utilize this knowledge is by fine-tuning the model. However, fine-tuning may require non-trivial computational resources and thus, previous works have focused on developing algorithms for conditional generation using only the pre-trained model (Chung et al., 2023; Rout et al., 2023; Chung et al., 2024; Yu et al., 2023; He et al., 2024). These methods modify the diffusion sampling process by computing additional gradient terms that move the sample towards the condition while denoising. When these gradients are computed using backpropagation through the model weights, there is a significant increase in inference time. On the other hand, when attempting to save computation by not propagating the condition information through the model, the generated image fails to capture the necessary long-range correlations.
As an example, in Figure 1, we use different algorithms to fill in the missing half of an image. Methods that backpropagate through the denoiser model (LDPS, PSLD (Rout et al., 2023), FreeDoM (Yu et al., 2023)) require significantly more time to run and do not consistently produce realistic results. Algorithms that do not compute gradients through the denoiser (MPGD (He et al., 2024)) fail to propagate the condition to distant pixels.
With the shortcomings of existing approaches in mind, we pose the following question: Can we avoid backpropagation through the model weights while at the same time maintaining long-range consistency when updating the sample with the conditioning signal? We observe that when a constraint increases the intensity of pixels in the predicted clean image, these pixel values are also expected to increase in the noisy image. From the diffusion point of view, the diffusion latent ${\bm{x}}_{t-1}$ is a linear combination of the current noisy image ${\bm{x}}_{t}$ and the predicted final image $\hat{{\bm{x}}}_{0}$ (Song et al., 2020), therefore any change in ${\bm{x}}_{t}$ can be also performed by changes in $\hat{{\bm{x}}}_{0}$.
This observation tells us that changes computed in the noisy and clean images can be interchanged. We find that this can be exploited to devise a new training-free conditional sampling algorithm that employs a numerical approximation for the expensive gradient computations. This allows us to speed up inference time while maintaining long-range dependencies in the image (Figure1). We take inspiration from Newton’s method and pose the proposed algorithm as an approximation to Newton steps that avoids computing the Jacobian inverse by using the Jacobian matrix directly.
We test the effectiveness of our proposed approach by generating images under both linear and non-linear constraints. We first show that our approach matches the results of state-of-the-art methods on free-form inpainting and 8×8$\times$8 × super-resolution at a fraction of the inference time. We then demonstrate how existing methods fail at inpainting large regions (25% of the image), while our algorithm obtains results closer to explicitly fine-tuning the diffusion model on inpainting. Finally, we apply our algorithm on non-linear constraints to perform style-guided and mask-guided generation, where we show that the proposed method synthesizes images that satisfy the constraints better than existing approaches.
@@ 2.1 Denoising Diffusion Models
Denoising diffusion models, proposed for image generation by Ho et al. (2020), have now been widely adopted mainly due to their exceptional ability to synthesize diverse and high-quality samples. The original formulation views the training and inference process as a hierarchical latent variable model ${\bm{x}}_{T}\rightarrow{\bm{x}}_{T-1}\rightarrow\dots\rightarrow{\bm{x}}_{1}\rightarrow{\bm{x}}_{0}$, where the final latent is distributed normally ${\bm{x}}_{T}~{}\sim N(\mathbf{0},{\bm{I}})$ and $p({\bm{x}}_{0})$ represents the data distribution. After choosing a noise schedule $a_{t}$ that defines the forward transitions ${\bm{x}}_{t}\rightarrow{\bm{x}}_{t+1}$ that corrupt the data with Gaussian noise, usually centered at $\sqrt{\frac{a_{t+1}}{a_{t}}}{\bm{x}}_{t}$ and with variance $(1-\frac{a_{t+1}}{a_{t}})$, the model is trained to reverse each individual step in the diffusion process by predicting the noise added to the clean sample. In this formulation, the predicted noise is shown to approximate the score function $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ (Song et al., 2021) of the diffusion latent variables.
Further iterations of denoising diffusion introduced classifier guidance (Dhariwal & Nichol, 2021), which adapted a pre-trained unconditional diffusion model for conditional sampling by training an additional classifier $p({\bm{y}}|{\bm{x}}_{t})$. During inference, each reverse step is also perturbed by the gradient $\nabla_{{\bm{x}}_{t}}\log p({\bm{y}}|{\bm{x}}_{t})$ which guides the diffusion latent ${\bm{x}}_{t}$ towards regions that also satisfy the condition ${\bm{y}}$. Classifier-free guidance (Ho & Salimans, 2022) eliminated the need for training an additional classifier to perform conditional generation by incorporating it into the base diffusion model training process.
Latent diffusion models (LDMs) proposed Rombach et al. (2022), reduced the complexity of training large diffusion models by training the diffusion in the latent space of a pre-trained autoencoder. By utilizing an encoder $\mathcal{E}$ and decoder $\mathcal{D}$ that accurately reconstruct images ${\bm{x}}_{0}\approx\mathcal{D}(\mathcal{E}({\bm{x}}_{0}))$, the diffusion process can be made more efficient as the learned redundant information in an image is left for the decoder to reconstruct. Most large text-to-image diffusion models, such as Stable Diffusion (Podell et al., 2024), are based on this encoder-decoder approach to reduce complexity.
@@ 2.2 Gradient descent steps for constrained sampling
Previous works have studied whether large pre-trained diffusion models, which required significant investment to train, can be directly used for inference under novel conditions without additional tuning for each different constraint (Chung et al., 2023). The most typical problem formulation is denoising the sequence ${\bm{x}}_{T},{\bm{x}}_{T-1},...,{\bm{x}}_{1},{\bm{x}}_{0}$ under the linear constraint on the final signal in the form ${\bm{A}}{\bm{x}}_{0}={\bm{y}}$, or using a relaxed version that requires minimization of $||{\bm{A}}{\bm{x}}_{0}-{\bm{y}}||_{2}^{2}$, possibly as part of the likelihood function $p({\bm{y}}|{\bm{x}}_{0})={\cal N}({\bm{y}};{\bm{A}}{\bm{x}}_{0},\sigma^{2}{\bm{I}})$.
In contrast to classifier guidance, which explicitly trained a separate model for the likelihood $p({\bm{y}}\mid{\bm{x}}_{t})$, the linear constraint only applies to the final, noise-free image ${\bm{x}}_{0}$. Thus, existing methods rely on Tweedie’s formula (Efron, 2011), by which denoising diffusion models approximating $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ can be used to express the expected value of ${\bm{x}}_{0}$, denoted as $\hat{{\bm{x}}}_{0}$. Including the constraint as if an additional observed variable ${\bm{y}}$ was generated requires the addition of the term $\nabla_{{\bm{x}}_{t}}\log p({\bm{y}}\mid{\bm{x}}_{t})$ to every diffusion sampling step, and most previous methods focused on different approximations of $p({\bm{y}}\mid{\bm{x}}_{t})$ using the estimated $\hat{{\bm{x}}}_{0}$.
@@
For example, in the case of inpainting missing pixels, the matrix ${\bm{A}}$ extracts the subsection of the known pixels in image ${\bm{x}}_{0}$ to be compared with a given target ${\bm{y}}$. The estimated expected value of ${\bm{x}}_{0}$ at the end of the chain is provided by the diffusion model as a nonlinear function $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$ learned by the denoiser network during training. Typically, these moves are gradient descent moves, i.e. moves of ${\bm{x}}_{t}$ in the direction of $-{\bm{e}}_{t}$ where
$-{\bm{e}}_{t}=-\nabla_{{\bm{x}}_{t}}C({\bm{x}}_{t})=-{\bm{J}}^{T}{\bm{A}}^{T}({\bm{A}}\hat{{\bm{x}}}_{0}-{\bm{y}})=-{\bm{J}}^{T}{\bm{e}},$ (2)
${\bm{J}}=\nabla_{{\bm{x}}_{t}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t}),\quad{\bm{e}}={\bm{A}}^{T}({\bm{A}}\hat{{\bm{x}}}_{0}-{\bm{y}}).$ (3)
Note that we name the error signal ${\bm{e}}$ as the (negative) direction of the gradient w.r.t. ${\bm{x}}_{0}$ itself, and the ${\bm{e}}_{t}$ is the matching move in ${\bm{x}}_{t}$.
The matrix ${\bm{A}}^{T}$ inverses the operation of ${\bm{A}}$ and can usually be computed for each task a-priori. However, computation of the full Jacobian ${\bm{J}}$ is impractical, so the gradient is instead computed using backpropagation through $C({\bm{x}}_{t})$. In (Chung et al., 2023) and (Chung et al., 2024) for example, gradient steps of this form are applied to ${\bm{x}}_{t}$ at each step $t$ of generation before moving on to the next stage. As optimization of ${\bm{x}}_{t}$ might reduce the total noise in the image below what the denoising at $t-1$ was trained for, the gradient steps moving ${\bm{x}}_{t}$ towards optimizing $C({\bm{x}}_{t})$ can be combined with adding additional noise, which could also be seen as a form of stochastic averaging, as done by Yu et al. (2023).
3 Method
3.1 Approximate Newton steps
We start by observing that a Newton optimization step, instead of moving ${\bm{x}}_{t}$ in the gradient descent direction ${\bm{J}}^{T}{\bm{e}}$, moves it in the direction ${\bm{J}}^{-1}{\bm{e}}$. We demonstrate that on a more general form of the cost
@@ 3.1 Approximate Newton steps
for some target ${\bm{y}}$ to be matched with projection function $f$, which in the linear case is $f({\bm{x}})={\bm{A}}{\bm{x}}$. The Gauss-Newton optimization would first approximate
@@
where ${\bm{J}}_{f}$ is the Jacobian of $f$. We then rewrite the cost of move $-{\bm{e}}_{t}$ as
$C({\bm{x}}_{t}-{\bm{e}}_{t})=$ $(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{J}}_{f}{\bm{J}}{\bm{e}}_{t}-{\bm{y}})^{T}$
$(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{J}}_{f}{\bm{J}}{\bm{e}}_{t}-{\bm{y}}).$ (6)
@@
assuming the inverses exist. In the case of $f({\bm{x}})={\bm{A}}{\bm{x}}$ for inpainting and superresolution tasks, where ${\bm{y}}$ is lower-dimensional, the inverse of ${\bm{J}}_{f}$, obviously is not defined, but we can use the pseudoinverse. Similarly, for a non-linear $f$, we can compute ${\bm{e}}$ using numerical methods, e.g. backpropagation through $f$ when it is a neural network.
@@
Therefore, ${\bm{e}}$ is the same in both gradient descent and Newton optimization, but the relationship to ${\bm{e}}_{t}$ is different:
$\text{GD}:\ {\bm{e}}_{t}={\bm{J}}^{T}{\bm{e}},\ \ \text{Newton}:\ {\bm{J}}{\bm{e}}_{t}={\bm{e}}$ (9)
@@
The gradient descent step can be computed without directly evaluating the Jacobian by back-propagation on the scalar cost $C$. For the computation of ${\bm{J}}^{-1}{\bm{e}}$ on the other hand, we have no such method. However, we can efficiently compute the move
$\epsilon{\bm{e}}_{t}=\epsilon{\bm{J}}{\bm{e}}.$ (10)
We follow the Newton recipe suggesting multiplication of the error by ${\bm{J}}$ in Eq. (9) but multiply ${\bm{e}}$ by it instead of ${\bm{e}}_{t}$. Intuitively, this is justified by the fact that the mode of ${\bm{x}}_{t-1}$ in the next stage of denoising is a weighted sum of $\hat{{\bm{x}}_{0}}$ and ${\bm{x}}_{t}$, meaning that the changes in ${\bm{x}}_{t}$ can be equivalently represented by proportional moves in the same direction of either component. This is the consequence of the modeling assumption that in the posterior the stages are separated by independent Gaussian noise, and early inpaiting models simply assumed ${\bm{e}}_{t}={\bm{e}}$, missing on the effect of the image prior $p({\bm{x}}_{0})$, which is implicit in the learned function ${\bm{x}}_{0}({\bm{x}}_{t})$. In addition to that, in Appendix A.1 we provide an analysis of the convergence on the algorithm using inexact Newton methods (Dembo et al., 1982).
@@
The direction of optimization we propose in Eq. (10) has at least two advantages over the usual gradient descent update. First, the direction $\epsilon{\bm{e}}_{t}$ can be computed numerically to save both on computation (2-fold) and memory (2.5-fold) compared to using backpropagation on the cost in Eq. (1). To derive that update, consider the function $f(s)=\hat{{\bm{x}}}_{0}({\bm{x}}_{t}-s{\bm{e}})$ where the variable $s$ is scalar. Its derivative at $s=0$ is
$\frac{df}{ds}=-{\bm{J}}{\bm{e}},$ (11)
and so the direction ${\bm{h}}$ can be approximated numerically using the approximate derivative of $f(s)$ at $s=0$
$f^{\prime}=\frac{1}{\delta}[\hat{{\bm{x}}}_{0}({\bm{x}}_{t}+\delta{\bm{e}})-\hat{{\bm{x}}}_{0}({\bm{x}}_{t})],$ (12)
${\bm{e}}_{t}=-\epsilon{\bm{J}}{\bm{e}}\approx\frac{\epsilon}{\delta}[\hat{{\bm{x}}}_{0}({\bm{x}}_{t}+\delta{\bm{e}})-\hat{{\bm{x}}}_{0}({\bm{x}}_{t})],$ (13)
@@
requiring no backpropagation but instead two forward passes through the network: one to compute $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$ and another to compute $\hat{{\bm{x}}}_{0}$ for the ${\bm{x}}_{t}$ perturbed in the negative direction of the error vector.
@@
The second advantage comes from the asymmetry of the Jacobian ${\bm{J}}$. The theoretically optimal ${\bm{x}}_{0}({\bm{x}}_{t})$ would have a symmetric Jacobian, as $\hat{{\bm{x}}}_{0}$ approximates the true expectation $E[{\bm{x}}_{0}|{\bm{x}}_{t}]=\frac{1}{\sqrt{\alpha_{t}}}[{\bm{x}}_{t}+{\bm{v}}_{t}\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})]$, and the gradient of this is indeed symmetric:
@@
This would render the gradient descent and the proposed update equivalent. In Section 3.2, we show that in practice, the trained denoiser models do not satisfy this ideal condition, making the two update directions noticeably different. Whether the denoising diffusion models are trained with score matching in mind (Song et al., 2021) or using the variational method of Ho et al. (2020), they do not directly optimize to match the real score $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ everywhere nor are they constrained to produce symmetric Jacobians. We hypothesize that this discrepancy in the two update directions is one of the reasons why our optimization of ${\bm{x}}_{t}$ may be more suitable for some applications, as demonstrated in Section 4.

### v3

@@ Abstract
Large denoising diffusion models, such as Stable Diffusion, have been trained on billions of image-caption pairs to perform text-conditioned image generation. As a byproduct of this training, these models have acquired general knowledge about image statistics, which can be useful for other inference tasks. However, when confronted with sampling an image under new constraints, e.g. generating the missing parts of an image, using large pre-trained text-to-image diffusion models is inefficient and often unreliable. Previous approaches either utilized backpropagation through the denoiser network, making them significantly slower and more memory-demanding than simple text-to-image generation, or only enforced the constraint locally, failing to capture critical long-range correlations in the sampled image. In this work, we propose an algorithm that enables fast, high-quality generation under arbitrary constraints. We show that in denoising diffusion models, we can employ an approximation to Newton’s optimization method that allows us to speed up inference and avoid the expensive backpropagation operations. Our approach produces results that rival or surpass the state-of-the-art training-free inference methods while requiring a fraction of the time. We demonstrate the effectiveness of our algorithm under both linear (inpainting, super-resolution) and non-linear (style-guided generation) constraints. An implementation is provided at this GitHub repository.
@@ 1 Introduction
The development of large text-to-image models [22, 27, 26, 30] has made denoising diffusion [32, 16] the go-to approach for capturing complex data distributions in high-dimensional spaces, such as images. By training on billions of text-image pairs, these models have acquired general knowledge about the image space, beyond text-to-image generation. This knowledge is useful in quickly adapting to new conditions [45, 40] and utilizing model features to solve downstream image tasks [38, 39].
The simplest way to utilize this prior knowledge is by fine-tuning the model. However, fine-tuning may require non-trivial computational resources and thus, previous works have focused on developing algorithms for conditional generation using only the pre-trained model [5, 28, 6, 44, 13]. These methods modify the diffusion sampling process by computing additional gradient terms that move the sample towards the condition while denoising. When these gradients are computed using backpropagation through the model weights, there is a significant increase in inference time. On the other hand, when attempting to save computation by not propagating the condition information through the model, the generated image fails to capture the necessary long-range correlations.
As an example, in Figure 1, we use different algorithms to fill in the missing half of an image. Methods that backpropagate through the denoiser model (LDPS, PSLD [28], FreeDoM [44]) require significantly more time to run and do not consistently produce realistic results. Algorithms that do not compute gradients through the denoiser (MPGD [13]) fail to propagate the condition to distant pixels. With the shortcomings of existing approaches in mind, we pose the following question: Can we avoid backpropagation through the model weights while at the same time maintaining long-range consistency when updating the sample with the conditioning signal?
We view the denoiser network as a function that removes all noise from an input image. In that context, the goal of constrained sampling is to find how the input should change such that the clean output satisfies the condition better. This involves the denoising function’s Jacobian matrix, which transforms a local change in the input to a change in outputs. We show that existing methods that backpropagate the constraint through the model employ the transpose of the Jacobian to invert this transformation, corresponding to the gradient descent direction.
An alternative direction is given from the Newton method [24], but requires more computation as it involves finding the Jacobian inverse. However, the input and output of the denoiser are closely related; when the intensity of some pixels in the final, noise-free image is increased, the same pixels are also expected to brighten in the noisy image and vice versa. Therefore, we propose to approximate the Jacobian inverse in the Newton step with the Jacobian matrix itself. We show that this is a valid approximation that (a) is fundamentally different from the gradient descent direction used in previous works, (b) is cheap to compute, with only two forward passes through the denoiser required, and (c) can quickly converge to the desired solution in large pre-trained diffusion models.
To evaluate, we generate images under both linear and non-linear constraints. We first show that our approach matches the results of state-of-the-art methods on free-form inpainting and $8\times$ super-resolution at a fraction of the inference time. We then demonstrate how existing methods fail at inpainting large regions, while our algorithm obtains results closer to a fully fine-tuned diffusion model on inpainting. Finally, we show how we can apply our algorithm to non-linear constraints and perform style-guided and mask-guided generation, where the proposed method consistently generates images that satisfy the constraints better than existing approaches.
@@
[28] [44] [13]
Figure 1: When tasked with completing the missing half of an image, previous methods are slow and fail to capture the important long-range dependencies between pixels. The proposed algorithm generates a reasonable image at a fraction of the time.
@@ 2.1 Denoising Diffusion Models
Denoising diffusion models [32, 16] have been widely adopted due to their exceptional ability to synthesize diverse and high-quality samples. The original formulation treats the training and inference process as a hierarchical latent variable model ${\bm{x}}_{T}\rightarrow{\bm{x}}_{T-1}\rightarrow\dots\rightarrow{\bm{x}}_{1}\rightarrow{\bm{x}}_{0}$, where the final latent is distributed normally ${\bm{x}}_{T}~\sim N(\mathbf{0},{\bm{I}})$ and $p({\bm{x}}_{0})$ represents the data distribution. Given a noise schedule $\alpha_{t}$ that defines the forward transitions ${\bm{x}}_{t}\rightarrow{\bm{x}}_{t+1}$ that corrupt the data with Gaussian noise, usually centered at $\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}{\bm{x}}_{t}$ and with variance $(1-\frac{\alpha_{t+1}}{\alpha_{t}})$, the model is trained to reverse each step in the diffusion process by predicting the noise added to the clean sample. The predicted noise is shown to approximate the score function $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ [34] of the diffusion latent variables.
Further iterations of denoising diffusion introduced classifier guidance [9], which adapted a pre-trained unconditional diffusion model for conditional sampling by training an additional classifier $p({\bm{y}}|{\bm{x}}_{t})$. During inference, each reverse step also includes the gradient $\nabla_{{\bm{x}}_{t}}\log p({\bm{y}}|{\bm{x}}_{t})$, which guides the diffusion latent ${\bm{x}}_{t}$ towards regions that also satisfy the condition $\bm{y}$}. However, if the conditioning $\bm{y}$} is already given, training this additional classifier on top of the diffusion model is costly. Classifier-free guidance [15] eliminated the need for an additional classifier by incorporating the conditional guidance into the base diffusion model training process.
The most widely adopted formulation of denoising diffusion is Latent diffusion models (LDMs) [27]. LDMs reduce the complexity by modeling the compressed latent space of an autoencoder instead of images. With an encoder $\mathcal{E}$ and decoder $\mathcal{D}$ that accurately reconstruct images ${\bm{x}}_{0}\approx\mathcal{D}(\mathcal{E}({\bm{x}}_{0}))$, the diffusion process can be made more efficient as redundant information in an image is left for the decoder to reconstruct. Most large text-to-image diffusion models, such as Stable Diffusion [27, 23], are based on the LDM approach.
@@ 2.2 Gradient descent steps for constrained sampling
Previous works studied whether large pre-trained diffusion models, which required significant investment to train, can be directly used for inference under novel conditions without additional tuning for each different constraint [5]. The typical problem formulation is denoising the sequence ${\bm{x}}_{T},{\bm{x}}_{T-1},...,{\bm{x}}_{1},{\bm{x}}_{0}$ under a linear constraint on the final signal in the form ${\bm{A}}{\bm{x}}_{0}={\bm{y}}$, or a relaxed version that minimizes ${\bm{A}}{\bm{x}}_{0}-{\bm{y}}||_{2}^{2}$ as part of the likelihood $p({\bm{y}}|{\bm{x}}_{0})={\cal N}({\bm{y}};{\bm{A}}{\bm{x}}_{0},\sigma^{2}{\bm{I}})$.
Contrary to classifier guidance, which trained a separate model for the likelihood $p({\bm{y}}\mid{\bm{x}}_{t})$, the linear constraint only applies to the final, noise-free image ${\bm{x}}_{0}$. Thus, existing methods rely on Tweedie’s formula [10], by which denoising diffusion models approximating $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ can be used to express the expected value of ${\bm{x}}_{0}$, denoted as $\hat{{\bm{x}}}_{0}$. Including the constraint as if an additional observed variable $\bm{y}$} was generated requires adding $\nabla_{{\bm{x}}_{t}}\log p({\bm{y}}\mid{\bm{x}}_{t})$ to every diffusion sampling step, and previous works considered different approximations of $p({\bm{y}}\mid{\bm{x}}_{t})$ using the estimated $\hat{{\bm{x}}}_{0}$ [5, 44].
@@
For example, in the case of inpainting missing pixels, the matrix $\bm{A}$} extracts the subsection of the known pixels in image $\hat{{\bm{x}}}_{0}$ to be compared with a given target $\bm{y}$}. The estimated expected value of ${\bm{x}}_{0}$ at the end of the chain is provided by the diffusion model as a nonlinear function $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$ learned by the denoiser network during training. Typically, these moves are gradient descent moves, i.e. moves of ${\bm{x}}_{t}$ in the direction of ${\bm{e}}_{t}$ where
${\bm{e}}_{t}=\nabla_{{\bm{x}}_{t}}C({\bm{x}}_{t})={\bm{J}}^{T}{\bm{A}}^{T}({\bm{A}}\hat{{\bm{x}}}_{0}-{\bm{y}})={\bm{J}}^{T}{\bm{e}},$ (2)
${\bm{J}}=\nabla_{{\bm{x}}_{t}}\hat{{\bm{x}}}_{0}({\bm{x}}_{t}),\quad{\bm{e}}={\bm{A}}^{T}({\bm{A}}\hat{{\bm{x}}}_{0}-{\bm{y}}).$
Note that we name the error signal $\bm{e}$} as the (negative) direction of the gradient w.r.t. ${\bm{x}}_{0}$ itself, and the ${\bm{e}}_{t}$ is the matching move in the noisy ${\bm{x}}_{t}$.
The matrix ${\bm{A}}^{T}$ inverses the operation of $\bm{A}$} and can usually be computed for each task a-priori. Since computing the full Jacobian $\bm{J}$} is impractical, the gradient is instead computed using backpropagation through $C({\bm{x}}_{t})$, which yields the direction ${\bm{J}}^{T}{\bm{e}}$. Chung et al. [5], [6], for example, apply gradient steps of this form to ${\bm{x}}_{t}$ at each step $t$ of generation before moving on to the next stage. As optimization of ${\bm{x}}_{t}$ might reduce the total noise in the image below what the denoising at t1t-1 was trained for, the gradient steps moving ${\bm{x}}_{t}$ towards optimizing $C({\bm{x}}_{t})$ can be combined with adding additional noise, which could also be seen as a form of stochastic averaging, as done by Yu et al. [44].
3 Method: Approximate Newton steps
We start by observing that a Newton optimization step, instead of moving ${\bm{x}}_{t}$ in the gradient descent direction ${\bm{J}}^{T}{\bm{e}}$, moves it in the direction ${\bm{J}}^{-1}{\bm{e}}$. We demonstrate it on a more general form of the cost
@@ 3 Method: Approximate Newton steps
for some target $\bm{y}$} to be matched with projection function $f$, which in the linear case is $f({\bm{x}})={\bm{A}}{\bm{x}}$. The Newton optimization would first approximate
@@
where ${\bm{J}}_{f}$ is the Jacobian of $f$. We then rewrite the cost of move ${\bm{e}}_{t}$ as
$C({\bm{x}}_{t}-{\bm{e}}_{t})=(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{J}}_{f}{\bm{J}}{\bm{e}}_{t}-{\bm{y}})^{T}(f(\hat{{\bm{x}}}_{0}({\bm{x}}_{t}))-{\bm{J}}_{f}{\bm{J}}{\bm{e}}_{t}-{\bm{y}}).$ (5)
@@
assuming the inverses exist. In the case of $f({\bm{x}})={\bm{A}}{\bm{x}}$ for inpainting and super-resolution tasks, where $\bm{y}$} is lower-dimensional, the inverse of ${\bm{J}}_{f}$, is not defined, but we can use the pseudo-inverse. Similarly, for a non-linear $f$, we can compute $\bm{e}$} using numerical methods, e.g. backpropagation through $f$ when it is a neural network, which would approximate $\bm{e}$} using ${\bm{J}}_{f}^{T}$.
@@
Therefore, $\bm{e}$} is the same in gradient descent and Newton optimization, but its relationship to ${\bm{e}}_{t}$ differs
$\text{(a)}\ \text{GD}:\ {\bm{e}}_{t}={\bm{J}}^{T}{\bm{e}},\quad\text{(b)}\ \text{Newton}:\ {\bm{J}}{\bm{e}}_{t}={\bm{e}}.$ (8)
@@
Gradient descent can be computed without directly evaluating the Jacobian by backpropagation on the scalar cost $C$. For the computation of ${\bm{J}}^{-1}{\bm{e}}$, on the other hand, we have no such method.
In cases where computing the inverse of the Jacobian is prohibitive, inexact Newton methods [7] propose first finding an approximation ${\bm{e}}_{t}^{*}$ to the solution of Eq (8) (b), performing the Newton step using the approximate solution, and reiterating until convergence. When the residual $\bm{r}$} is strictly reduced at every step, inexact Newton methods converge to the correct solution
Since we know that the input and output of the denoiser are related by additive Gaussian noise, we propose the inexact move
${\bm{e}}_{t}={\bm{J}}{\bm{e}}$ (10)
where instead of computing the Jacobian inverse, we use the Jacobian-error vector product. If the Jacobian tells us how to transform a local change in the input to a change in the output, then we posit that, for denoising diffusion models, we can use the same transformation for the inverse. Substituting ${\bm{e}}_{t}^{*}$ with the proposed update, adding a learning rate $\lambda$ to control convergence, we get the residual
${\bm{r}}={\bm{e}}-\lambda{\bm{J}}^{2}{\bm{e}}=({\bm{I}}-\lambda{\bm{J}}^{2}){\bm{e}}.$ (11)
For this residual to be strictly reduced at every iteration of our algorithm, we require
As long as the spectrum of the denoiser Jacobian matrix is bounded, we can choose a small enough $\lambda$ that strictly reduces the residual. However, the proposed move would converge slowly if the only $\lambda$ allowed were very small. In practice, we find that we can use large learning rates, making the convergence similar or even better than gradient descent. In Appendix A.1, we present an analysis of the spectral properties of the Jacobian matrix of the Stable Diffusion denoiser used in our experiments. Furthermore, in Appendix A.3 we demonstrate the differences between our proposed inexact and the exact Newton method on a small-scale experiment.
Figure 2: We showcase the difference between the proposed method to compute the update direction ($\bm{J}$}) and gradient descent (${\bm{J}}^{T}$). The heatmaps indicate where the input ${\bm{x}}_{t}$ would change when perturbing a single pixel in the output, denoted in green. The two directions are considerably different, with ours capturing better longer-range correlations and maintaining shapes. Even though we use finite differences, the direction computed from $\bm{J}$} is sharper in some regions, like the outlines.
@@
The direction of optimization we propose in Eq. (10) has another advantage over the gradient descent update. The direction ${\bm{e}}_{t}$ can be computed numerically to save both on computation and memory compared to using backpropagation on the cost in Eq. (1). To derive the update, consider the function $h(s)=\hat{{\bm{x}}}_{0}({\bm{x}}_{t}+s{\bm{e}})$ where the variable $s$ is scalar. Its derivative at $s=0$ is
$\left.\frac{dh}{ds}\right\rvert_{s=0}={\bm{J}}{\bm{e}},$ (13)
and so the direction ${\bm{e}}_{t}$ can be computed using the numerical derivative of $h(s)$ at $s=0$
${\bm{e}}_{t}=\lambda{\bm{J}}{\bm{e}}=\lambda\left.\frac{dh}{ds}\right\rvert_{s=0}\approx\lambda\frac{h(\delta)-h(0)}{\delta}=\lambda\frac{\hat{{\bm{x}}}_{0}({\bm{x}}_{t}+\delta{\bm{e}})-\hat{{\bm{x}}}_{0}({\bm{x}}_{t})}{\delta},$ (14)
@@
requiring no backpropagation but instead two forward passes through the network: one to compute $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$ and another to compute $\hat{{\bm{x}}}_{0}$ for the ${\bm{x}}_{t}$ perturbed in the direction of the error vector.
@@
It is important to point out that the theoretically optimal ${\bm{x}}_{0}({\bm{x}}_{t})$ would have a symmetric Jacobian, as $\hat{{\bm{x}}}_{0}$ approximates the true expectation $E[{\bm{x}}_{0}|{\bm{x}}_{t}]=\frac{1}{\sqrt{\alpha_{t}}}[{\bm{x}}_{t}+{\bm{v}}_{t}\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})]$, and the gradient of this is indeed symmetric
@@
This would render the gradient descent and the proposed update equivalent. Yet, we find that the trained denoiser models do not satisfy this ideal condition, making the two update directions noticeably different.
@@
In Figure 2, we visualize this difference between $\bm{J}$} and ${\bm{J}}^{T}$ by computing rows of $\bm{J}$} and ${\bm{J}}^{T}$, and averaging the magnitude of all matrix values that contribute to each pixel. This indicates that there is a difference in which pixels in ${\bm{x}}_{t}$ would change when perturbing a single pixel in $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$ for the two update directions. Further experiments on the difference between the two update directions are included in Appendix A.2. Whether the denoising diffusion models are trained with score matching in mind [34] or using the variational method of Ho et al. [16], they do not directly optimize to match the real score $\nabla_{{\bm{x}}_{t}}\log p_{t}({\bm{x}}_{t})$ everywhere nor are they constrained to produce symmetric Jacobians. We hypothesize that this discrepancy in the two update directions is another reason why our optimization of ${\bm{x}}_{t}$ may be more suitable for some applications, which we demonstrate in the experiments (Section 4).